OpenStudy (anonymous):

solve each quadratic equation if the solutions are imaginary, write in a(plus or minus_+)bi form 5x^2-6x+3=0 8x^2-5x+2=0 4x^2+20x+28=0

OpenStudy (jdoe0001):

how are you supposed to solve them? using the quadratic formula, or factoring, or "completing the square" or ?

OpenStudy (anonymous):

using the quadratic formula

OpenStudy (jdoe0001):

then those are simple just plug and chug \(\large { \textit{quadratic formula}\\ \begin{array}{llll} {\color{blue}{ 5}}x^2{\color{red}{ -6}}x{\color{green}{ +3}}=0\\ {\color{blue}{ 8}}x^2{\color{red}{ -5}}x{\color{green}{ +2}}=0\\ {\color{blue}{ 4}}x^2{\color{red}{ +20}}x{\color{green}{ +28}}=0 \end{array} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}} }\)

OpenStudy (anonymous):

\[6\pm \sqrt{-24}\over 10\] and thats as far as ive gotten for the first one

OpenStudy (anonymous):

where do you go from there